OpenStudy (anonymous):

Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 72 when x = 6.

OpenStudy (anonymous):

A. y = 12x B. y = 4x2 C. y = x2 + 25 D. y = 2x2

OpenStudy (wolfe8):

Varies directly means x=ky where k is some constant. So you plug in x and y to find k, then put k in.

OpenStudy (anonymous):

6=k*72?

OpenStudy (wolfe8):

Good call. I didn't see the square

jigglypuff314 (jigglypuff314):

"y varies directly with x2" means y = k x^2 (where k is the number you want to find) they gave you x and y, plug in so 72 = k (6^2) and solve for k

OpenStudy (anonymous):

Yea, so k=2 right?

jigglypuff314 (jigglypuff314):

um yeah :)

jigglypuff314 (jigglypuff314):

so plug what you got for k back into y = k x^2

jigglypuff314 (jigglypuff314):

@wolfe8 you confused meh dere >,< I only just learned it kinda a month ago direct is y = kx not x = ky

OpenStudy (anonymous):

72=2*36?

jigglypuff314 (jigglypuff314):

mmm the plugging x and y part was temporary just plug in k = 2 into y = (k) x^2

OpenStudy (wolfe8):

Lol sorry. I wasn't in math mode. I'm talking about coffee and tights.

jigglypuff314 (jigglypuff314):

0.o lol what ... nvm, don't wanna know... 0.0

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

y=(2)x^2?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

but that is not even a answer choice lol

jigglypuff314 (jigglypuff314):

yeah it is

OpenStudy (anonymous):

Oh yea it is D :P

jigglypuff314 (jigglypuff314):

yup :P

OpenStudy (anonymous):

Ty